# Near, far: Patch-ordering enhances vision foundation models’ scene understanding

### v2

@@ Abstract
We introduce NeCo: Patch Neighbor Consistency, a novel self-supervised training loss that enforces patch-level nearest neighbor consistency across a student and teacher model. Compared to contrastive approaches that only yield binary learning signals, i.e. ‘attract’ and ‘repel’, this approach benefits from the more fine-grained learning signal of sorting spatially dense features relative to reference patches. Our method leverages differentiable sorting applied on top of pretrained representations, such as DINOv2-registers to bootstrap the learning signal and further improve upon them. This dense post-pretraining leads to superior performance across various models and datasets, despite requiring only 19 hours on a single GPU. This method generates high-quality dense feature encoders and establishes several new state-of-the-art results such as +5.5 % and +6% for non-parametric in-context semantic segmentation on ADE20k and Pascal VOC, +7.2% and +5.7% for linear segmentation evaluations on COCO-Things and -Stuff and improvements in the 3D understanding of multi-view consistency on SPair-71k, by more than 1.5%.
@@ 3 Patch Neighbor Consistency
To ensure robust nearest neighbor consistency, we compute the cross-entropy loss in both directions and sum the losses. The final training loss, incorporating both directions, for all the patches is:
$\mathcal{L}_{\texttt{NeCo}}=\sum_{i=1}^{N^{\prime}}\mathcal{L}_{\text{CE}}(Q^{t}_{i},Q^{s}_{i})+\mathcal{L}_{\text{CE}}(Q^{s}_{i},Q^{t}_{i})$
This ensures, in a differentiable manner, that the order of nearest neighbors is consistent between the student and teacher features in both directions.
@@ 4.2 Comparison to State-of-the-Art
Figure 2: In-context scene understanding benchmark. Dense nearest neighbor retrieval performance is reported across various training data proportions on two scene-centric datasets, Pascal VOC and ADE20k. The retrieved cluster maps are compared with the ground truth using Hungarian matching (Kuang et al., 2021), and their mIoU score is reported. For all models, ViT-S16 is used except for DINOv2R and NeCo, where it is ViT-S14. For full tables, refer to Appendix B.3
In this section, we first compare the quality of frozen features learned through NeCo  with state-of-the-art methods in in-context learning via nearest neighbor retrieval and unsupervised semantic segmentation tasks. Next, we demonstrate NeCo’s versatility by applying it to five different pretraining models and show it improves their dense features consistently. We then evaluate the transferability of our learned dense representations to other datasets by using linear head semantic segmentation and end-to-end fine-tuning with Segmentor (Strudel et al., 2021). Finally, we show that NeCo  improves the 3D understanding of different vision models, as proposed in  El Banani et al. (2024), using the multiview consistency experiment.
Visual In-Context Learning Evaluation. We compare our approach to a recently proposed benchmark (Balazevic et al., 2023) that evaluates in-context reasoning in vision models. Unlike traditional linear segmentation methods, this evaluation does not require fine-tuning or end-to-end training. Instead, it creates validation segmentation maps by matching patch-level, feature-wise nearest neighbor similarities between validation images (queries) and training samples (keys). This method, inspired by NLP strategies, tests how well models learn tasks from a few examples. The results are presented in Figure 2. As shown, NeCo outperforms prior state-of-the-art methods such as CrIBo and DINOv2R by 4% to 13% on Pascal and ADE20k across different fractions. The performance gap between NeCo and others increases, particularly in the data-efficiency regime. This improvement is due to our method’s explicit enforcement of patch-level nearest neighbor consistency, resulting in higher-quality patch-level representations that remain effective even with fewer images. In contrast, other methods that promote image-level (Balazevic et al., 2023) or object-level (Lebailly et al., 2023) consistency force consistency between a pooled vector of patches, potentially leading to inadequate semantic patch-level representations, particularly in smaller datasets. Our method’s ability to perform well with few images brings vision models one step closer to in-context learning style generalist reasoning. For complete tables and details on scaling the results to larger models, please refer to Appendix B.3. For visualizations please refer to Appendix C.
@@
Linear Semantic Segmentation Evaluation. In this experiment, we keep the pretrained backbone frozen and train a linear layer on top of the spatial features to solve a supervised semantic segmentation task. Bilinear interpolation is used to match the spatial feature resolution to the image size, enabling the application of pixel-wise cross-entropy loss. This setup provides a better evaluation of the pretrained models compared to end-to-end finetuning, where all learned parameters are overwritten. The results, reported in Table 2, show that NeCo surpasses CrIBo on all datasets by at least 10% and outperforms DINOv2R by 5% to 7%. These significant improvements demonstrate that patches representing the same object or object part have higher similarities in feature space compared to other methods, as a simple linear layer can utilize these features for strong semantic segmentation.
@@
End-to-End Full-Finetuning Evaluation. One advantage of self-supervised pretraining is the ability to transfer learned general semantic features to specialized downstream tasks, improving performance in an end-to-end finetuning setup. We evaluate this capability of NeCo  by adding a transformer-based decoder from Segmenter (Strudel et al., 2021) on top of the feature extractor and finetuning the entire network for semantic segmentation. The backbone’s spatial features are fed into a transformer decoder along with $K$ learnable class tokens. These class and spatial tokens are projected onto each other to obtain patch-level predictions, which are then upsampled to match the input image size, enforcing pixel-wise cross-entropy loss. We report the mIoU scores achieved on Pascal VOC, Pascal Context, COCO-Stuff, and ADE20k in Table 4. Despite all parameters being adapted, the results show that NeCo  learns superior features, leading to better performance in downstream tasks, outperforming CrIBo by approximately 4%. Notably, while DINOv2R has demonstrated strong transfer results in various tasks, including semantic segmentation, due to being trained on a massive dataset of 142M images and using a combination of dense and classification losses, NeCo  surpasses even this. By training only 19 GPU-hours on COCO, which is a fraction of the original compute, we obtain consistent gains, setting a new state-of-the-art.
@@
DINO-B16 26.7 14.9 35.4 15.6 20.6 19.3 18.1 33.7 11.2 19.4 23.9 16.3 16.3 18.8 11.1 12.9 29.5 10.1 19.65
+ NeCo 26.0 17.5 36.7 16.9 22.0 23.5 18.0 33.5 10.6 21.0 21.6 20.2 20.6 19.3 13.7 13.9 36.3 12.5 21.32
DINOV2R-B14 31.2 34.4 56.7 23.2 26.1 29.4 37.5 51.3 20.8 36.6 36.7 31.7 26.2 29.4 15.4 26.5 39.2 11.9 32.50
+NeCo 40.6 25.7 58.0 20.4 37.1 33.3 37.2 56.2 24.6 37.7 40.8 33.1 23.7 35.1 22.8 23.9 45.4 22.4 34.34
@@ A.1 Dense Post-Pretraining
Our Data augmentations are the same as in Ziegler & Asano (2022). More specifically, we use: random color-jitter, Gaussian blur, grayscale and multi-crop augmentations. Similarly, the global crop’s resolution is 224x224 and the local crop’s resolution is 96x96, for the all the experiments except when working with Dinov2 where we use 98x98 for local crops. Furthermore, our generated global and local crops have the constraint that they intersect at least by $1\%$ of the original image size.
@@ B.2 Full Clustering Tables
NeCo ViT-S/14 21M 17.8 69.4 72.6 18.2 61.2 64.5
@@
NeCo ViT-B/14 85M 18.6 64.2 71.8 13.3 61.3 65.5
@@ B.3 Full Visual In-Context Learning Tables
DINOv2R ViT-S/14 21M 19.6 22.8 30.1 35.9 53.0 57.9 68.2 75.0
NeCo ViT-S/14 21M 23.7 27.2 34.0 39.8 66.5 70.3 76.3 80.2
@@
DINOv2R ViT-B/14 85M 22.1 25.8 33.2 38.7 51.8 58.9 70.6 77.3
NeCo ViT-B/14 85M 29.1 33.7 39.2 44.2 67.0 71.4 77.5 83.5
@@ B.4 Object detection with ViTDet
We report the performance of ViTDet (Li et al., 2022) on COCO dataset for ViT-S with DINO and DINOv2R backbones and compare them with the NeCo  finetuned versions. Our method improves DINO and DINOv2R backbones, specially DINO, increasing the validation Box AP slightly from 42.9 to 43.0 and the Mask AP from 38.6 to 38.7.
@@ B.6 Generalizing NeCo  to broader tasks and datasets
we report the performance of various backbones, including DINOv2R finetuned with the NeCo loss, on out-of-distribution datasets such as Cityscapes (Cordts et al., 2016)(semantic segmentation) and NYUd (Couprie et al., 2013)(monocular depth estimation). As demonstrated in 13(a) and 13(c), NeCo  consistently enhances the generalization of all backbones on the Cityscapes dataset. Furthermore, even for tasks that diverge from semantic segmentation, such as depth estimation, NeCo  reduces the DINOv2R error by around 2%. These findings highlight that NeCo not only maintains but improves the generalization capability of DINOv2R features across diverse tasks. For completeness, we also include results on (Markus Gerke, 2014), showing that NeCo consistently enhances the performance, even on out-of-distribution datasets.
@@
DINO-S/16 40.7 43.7
iBot-S/16 43.0 44.8
TimeT-S/16 42.2 42.8
Leopard-S/16 44.5 45.5
Dinov2R-S/14 48.9 52.9
@@
ViT-B Dinov2 0.412
@@ B.7 Extra Ablations
In 14(e), we vary the sorting steepness, denoted by $\beta$, for both teacher and student networks to evaluate the influence of hard or soft nearest neighbor assignments. The performance improves when the teacher’s steepness is higher or equal to the student’s, consistent with previous findings (Caron et al., 2021). Our best results are achieved when both networks have equal steepness. However, extreme steepness values (e.g., 1000) harm performance. This is because sorting patch similarities lacks clear boundaries, and formulating it as a hard assignment can force incorrect orderings, negatively impacting performance.
@@
We show the performance across different training epochs in 14(f). As the table shows, even after just one epoch of training, DINOv2R improves by 1% to 3% across various metrics. The performance continues to increase with more training epochs, but the improvements become smaller after 25 epochs, which is the number used in the paper.
@@
We show the effect of using different metrics for computing pair-wise patch similarity in 14(c). As the table shows, Cosine similarity is consistently better than Euclidean distance. We have used cosine similarity in all our experiments.
@@
ROI-Align component can be removed if the multi-crop augmentation is omitted or by using the same Global Crop across the teacher and student branches. We ablate the effect of removing ROI align in 14(d). As shown by the results, this removal negatively affects performance due to using weaker self-supervised supervision through the augmentations.

### v3

@@ Abstract
We introduce NeCo: Patch Neighbor Consistency, a novel self-supervised training loss that enforces patch-level nearest neighbor consistency across a student and teacher model. Compared to contrastive approaches that only yield binary learning signals, i.e. ‘attract’ and ‘repel’, this approach benefits from the more fine-grained learning signal of sorting spatially dense features relative to reference patches. Our method leverages differentiable sorting applied on top of pretrained representations, such as DINOv2-registers to bootstrap the learning signal and further improve upon them. This dense post-pretraining leads to superior performance across various models and datasets, despite requiring only 19 hours on a single GPU. This method generates high-quality dense feature encoders and establishes several new state-of-the-art results such as +2.3 % and +4.2% for non-parametric in-context semantic segmentation on ADE20k and Pascal VOC, +1.6% and +4.8% for linear segmentation evaluations on COCO-Things and -Stuff and improvements in the 3D understanding of multi-view consistency on SPair-71k, by more than 1.5%.
@@ 3 Patch Neighbor Consistency
And the final training loss is defined as the sum of per-patch losses computed over all samples:
$\mathcal{L}_{\texttt{NeCo}}=\sum_{i=1}^{N^{\prime}}\mathcal{L}_{\text{CE}}(Q^{t}_{i},Q^{s}_{i})$
This ensures, in a differentiable manner, that the order of nearest neighbors is consistent between the student and teacher features.
@@ 4.2 Comparison to State-of-the-Art
Figure 2: In-context scene understanding benchmark. Dense nearest neighbor retrieval performance is reported across various training data proportions on two scene-centric datasets, Pascal VOC and ADE20k. The retrieved cluster maps are compared with the ground truth using Hungarian matching (Kuang et al., 2021), and their mIoU score is reported. For all models, ViT-B16 is used except for DINOv2R and NeCo, where it is ViT-B14. For full tables, refer to Appendix B.3
In this section, we first compare the quality of frozen features learned through NeCo  with state-of-the-art methods in in-context learning via nearest neighbor retrieval and unsupervised semantic segmentation tasks. Next, we demonstrate NeCo’s versatility by applying it to five different pretraining models and show it improves their dense features consistently. We then evaluate the transferability of our learned dense representations to other datasets by using linear head semantic segmentation and end-to-end fine-tuning with Segmentor (Strudel et al., 2021). Finally, we show that NeCo  improves the 3D understanding of different vision models, as proposed in  El Banani et al. (2024), using the multiview consistency experiment. We report the results of applying NeCo  to more backbones in Appendix B.
Visual In-Context Learning Evaluation. We compare our approach to a recently proposed benchmark (Balazevic et al., 2023) that evaluates in-context reasoning in vision models. Unlike traditional linear segmentation methods, this evaluation does not require fine-tuning or end-to-end training. Instead, it creates validation segmentation maps by matching patch-level, feature-wise nearest neighbor similarities between validation images (queries) and training samples (keys). This method, inspired by NLP strategies, tests how well models learn tasks from a few examples. The results are presented in Figure 2. As shown, NeCo outperforms prior state-of-the-art methods such as CrIBo and DINOv2R by 4% to 13% on Pascal and ADE20k across different fractions. The performance gap between NeCo and others increases, particularly in the data-efficiency regime. This improvement is due to our method’s explicit enforcement of patch-level nearest neighbor consistency, resulting in higher-quality patch-level representations that remain effective even with fewer images. In contrast, other methods that promote image-level (Balazevic et al., 2023) or object-level (Lebailly et al., 2023) consistency force consistency between a pooled vector of patches, potentially leading to inadequate semantic patch-level representations, particularly in smaller datasets. Our method’s ability to perform well with few images brings vision models one step closer to in-context learning style generalist reasoning. For complete tables and details of using other backbone variants, please refer to Appendix B.3. For visualizations please refer to Appendix C.
@@
Linear Semantic Segmentation Evaluation. In this experiment, we keep the pretrained backbone frozen and train a linear layer on top of the spatial features to solve a supervised semantic segmentation task. Bilinear interpolation is used to match the spatial feature resolution to the image size, enabling the application of pixel-wise cross-entropy loss. This setup provides a better evaluation of the pretrained models compared to end-to-end finetuning, where all learned parameters are overwritten. The results, reported in Table 2, show that NeCo surpasses CrIBo on all datasets by at least 10% and consistently outperforms DINOv2R, achieving gains of up to 4.8%. These significant improvements demonstrate that patches representing the same object or object part have higher similarities in feature space compared to other methods, as a simple linear layer can utilize these features for strong semantic segmentation.
@@
End-to-End Full-Finetuning Evaluation. One advantage of self-supervised pretraining is the ability to transfer learned general semantic features to specialized downstream tasks, improving performance in an end-to-end finetuning setup. We evaluate this capability of NeCo  by adding a transformer-based decoder from Segmenter (Strudel et al., 2021) on top of the feature extractor and finetuning the entire network for semantic segmentation. The backbone’s spatial features are fed into a transformer decoder along with $K$ learnable class tokens. These class and spatial tokens are projected onto each other to obtain patch-level predictions, which are then upsampled to match the input image size, enforcing pixel-wise cross-entropy loss. We report the mIoU scores achieved on Pascal VOC, Pascal Context, COCO-Stuff, and ADE20k in Table 4. Despite all parameters being adapted, the results show that NeCo  learns superior features, leading to better performance in downstream tasks, outperforming CrIBo by at least 3.4% across different datasets and backbones. Notably, while DINOv2R has demonstrated strong transfer results in various tasks, including semantic segmentation, due to being trained on a massive dataset of 142M images and using a combination of dense and classification losses, NeCo  surpasses even this. By training only 19 GPU-hours on COCO, which is a fraction of the original compute, we obtain consistent gains of up to 1.7%, setting a new state-of-the-art.
@@
DINO-B16 26.7 14.9 35.4 15.6 20.6 19.3 18.1 33.7 11.2 19.4 23.9 16.3 16.3 18.8 11.1 12.9 29.5 10.1 19.6
+ NeCo 26.0 17.5 36.7 16.9 22.0 23.5 18.0 33.5 10.6 21.0 21.6 20.2 20.6 19.3 13.7 13.9 36.3 12.5 21.3
DINOV2R-B14 31.2 34.4 56.7 23.2 26.1 29.4 37.5 51.3 20.8 36.6 36.7 31.7 26.2 29.4 15.4 26.5 39.2 11.9 32.5
+NeCo 40.6 25.7 58.0 20.4 37.1 33.3 37.2 56.2 24.6 37.7 40.8 33.1 23.7 35.1 22.8 23.9 45.4 22.4 34.3
@@ A.1 Dense Post-Pretraining
Our Data augmentations are the same as in Ziegler & Asano (2022). More specifically, we use: random color-jitter, Gaussian blur, grayscale and multi-crop augmentations. Similarly, the global crop’s resolution is 224x224 and the local crop’s resolution is 96x96, for the all the experiments except when working with Dinov2 where we use 518x518 for global crops and 98x98 for local crops. The only place where the resolution is different for NeCo using DINOv2 is in the ablation studies, where we use 224x224 for global crops and 98x98 for local crops. Furthermore, our generated global and local crops have the constraint that they intersect at least by $1\%$ of the original image size.
@@
Registers in Dinov2
We report results on DINOv2R (Darcet et al., 2024), DINOv2XR, and DINOv2 (Oquab et al., 2023). For DINOv2XR, we remove the registers to match the input patch structure of the original DINOv2 model. DINOv2XR serves as an interesting experiment to evaluate whether removing registers, thus simplifying the architecture to align with DINOv2, affects performance significantly. However, all our primary results and comparisons focus on DINOv2R.
@@
Table 7: Post-Training Configurations for Different Experimental Setups used in this work. This set of tables shows all the different Post-Training parameter configurations used for post-training models for various experiments. The last row of each configuration table shows the GPU hours that are approximately used for post-training the model with the respective configuration. The rest of the parameters are the same and are as explained in Appendix A.1. Note, that the for the configuration in 7(a), denotes that the model is not eventually trained for 25 epochs but instead stopped after around the GPU hours mentioned since the model does not provide significant improvements afterwards (at around the 7th epoch). Last DINOv2R-XR means DINOv2R but excluding registers from the checkpoint().
(a) DINOv2R Best
Param Value
Global Crop 518
Local Crop 98
Backbone LR $1\mathrm{e}{-6}$
Head LR $1\mathrm{e}{-4}$
Max Epochs $25^{*}$
Drop Registers No
GPU Hours 18.8
(b) DINOv2R-XR Best
Param Value
Global Crop 518
Local Crop 98
Backbone LR $1\mathrm{e}{-6}$
Head LR $1\mathrm{e}{-4}$
Max Epochs $25^{*}$
Drop Registers Yes
GPU Hours 18.8
(c) DINOv2 Best
Param Value
Global Crop 518
Local Crop 98
Backbone LR $1\mathrm{e}{-6}$
Head LR $1\mathrm{e}{-4}$
Max Epochs $25^{*}$
Drop Registers N/A
GPU Hours 18.8
(d) DINOv2R-XR Ablations
Param Value
Global Crop 224
Local Crop 98
Backbone LR $1\mathrm{e}{-6}$
Head LR $1\mathrm{e}{-4}$
Max Epochs $10$
Drop Registers Yes
GPU Hours 7.2
(e) All others than DINOv2R
Param Value
Global Crop 224
Local Crop 96
Backbone LR $1\mathrm{e}{-5}$
Head LR $1\mathrm{e}{-4}$
Max Epochs $25$
Drop Registers N/A
GPU Hours 17.5
Table 8: The Post-Training Configuration used for each experiment table.
Table Configuration
1(b) 7(a)
2 7(a)
3 7(e)
4 7(a)
 5 7(a) for DINOv2R
7(e) for DINO
Table Configuration
6(e) 7(d)
9 7(a)
10 7(a), 7(b), 7(c)
11 7(a), 7(b), 7(c)
 12 7(e) for DINO
7(a), 7(b), 7(c) for rest
Table Configuration
13 7(e)
14 7(e)
15(c) 7(a), 7(b), 7(c)
16(f) 7(d)
17 7(e)
@@ B.2 Full Clustering Tables
DINOv2 ViT-S/14 21M 17.1 53.5 58.2 19.7 51.6 53.8
DINOv2R ViT-S/14 21M 18.0 59.1 64.5 20.1 55.1 59.2
NeCo (DINOv2) ViT-S/14 21M 15.5 50.4 57.7 20.0 56.0 60.0
NeCo (DINOv2R) ViT-S/14 21M 18.5 66.5 68.9 22.8 59.7 63.8
NeCo(DINOv2XR) ViT-S/14 21M 17.8 69.4 72.6 18.2 61.2 64.5
@@
DINOv2 ViT-B/14 85M 15.5 52.8 56.9 22.6 54.0 54.9
DINOv2R ViT-B/14 85M 21.4 62.2 64.6 23.1 55.2 57.8
NeCo (DINOv2) ViT-B/14 85M 17.2 66.8 71.1 23.7 62.2 63.1
NeCo (DINOv2R) ViT-B/14 85M 17.2 72.2 71.9 17.3 62.1 64.6
NeCo (DINOv2XR) ViT-B/14 85M 18.6 64.2 71.8 13.3 61.3 65.5
@@ B.3 Full Visual In-Context Learning Tables
DINOv2XR ViT-S/14 21M 19.6 22.8 30.1 35.9 53.0 57.9 68.2 75.0
DINOv2 ViT-S/14 21M 22.4 25.8 33.6 38.9 55.7 61.8 72.4 77.0
DINOv2R ViT-S/14 21M 23.7 27.1 33.9 39.5 60.1 65.7 74.5 78.8
NeCo (DINOv2) ViT-S/14 21M 22.1 25.2 32.9 38.0 59.8 64.6 73.4 78.6
NeCo (DINOv2R) ViT-S/14 21M 23.7 27.2 34.7 40.9 65.6 70.1 76.8 80.7
NeCo (DINOv2XR) ViT-S/14 21M 23.7 27.2 34.0 39.8 66.5 70.3 76.3 80.2
@@
DINOv2XR ViT-B/14 85M 22.1 25.8 33.2 38.7 51.8 58.9 70.6 77.3
DINOv2 ViT-B/14 85M 22.1 25.2 32.3 37.9 54.1 60.5 71.5 76.7
DINOv2R ViT-B/14 85M 25.6 29.4 36.5 41.9 59.1 64.6 74.2 78.7
NeCo (DINOv2) ViT-B/14 85M 26.6 31.1 38.9 43.4 68.4 72.8 79.6 82.8
NeCo (DINOv2R) ViT-B/14 85M 27.8 32.1 39.7 44.5 69.0 73.1 79.8 82.9
NeCo (DINOv2XR) ViT-B/14 85M 29.1 33.7 39.2 44.2 67.0 71.4 77.5 83.5
@@ B.4 Object detection with ViTDet
We report the performance of ViTDet (Li et al., 2022) on COCO dataset for ViT-S with DINO, DINOv2, and DINOv2R backbones and compare them with the NeCo  finetuned versions. Our method consistently improves all the DINO family backbones, as shown by Table 12. In particular, NeCo improves validation Box AP and Mask AP by 2.4% and 3.3%, respectively, over DINOv2, highlighting the versatility of our approach
@@ B.6 Generalizing NeCo  to broader tasks and datasets
we report the performance of various backbones, including DINOv2R finetuned with the NeCo loss, on out-of-distribution datasets such as Cityscapes (Cordts et al., 2016)(semantic segmentation) and NYUd (Couprie et al., 2013)(monocular depth estimation). As demonstrated in 15(a) and 15(c), NeCo  consistently enhances the generalization of all backbones on the Cityscapes dataset. Furthermore, even for tasks that diverge from semantic segmentation, such as depth estimation, NeCo  reduces the DINOv2R error by around 2%. These findings highlight that NeCo not only maintains but improves the generalization capability of DINOv2R features across diverse tasks. For completeness, we also include results on (Markus Gerke, 2014) in 15(b), showing that NeCo consistently enhances the performance, even on out-of-distribution datasets.
@@
DINO-B/16 41.8 42.5
iBot-B/16 42.9 44.5
Leopard-B/16 43.8 44.5
DINOv2XR-B/14 52.3 56.0
DINOv2-B/14 51.4 57.7
DINOv2R-B/14 53.4 58.9
@@
DINOv2R 0.455
+ NeCo 0.455
ViT-B/14 DINOv2 0.412
+ NeCo 0.385
DINOv2XR 0.410
+ NeCo 0.397
@@ B.7 Extra Ablations
In 16(e), we vary the sorting steepness, denoted by $\beta$, for both teacher and student networks to evaluate the influence of hard or soft nearest neighbor assignments. The performance improves when the teacher’s steepness is higher or equal to the student’s, consistent with previous findings (Caron et al., 2021). Our best results are achieved when both networks have equal steepness. However, extreme steepness values (e.g., 1000) harm performance. This is because sorting patch similarities lacks clear boundaries, and formulating it as a hard assignment can force incorrect orderings, negatively impacting performance.
@@
We show the performance across different training epochs in 16(f). As the table shows, even after just one epoch of training, DINOv2R improves by 1% to 3% across various metrics. The performance continues to increase with more training epochs, but the improvements become smaller after 25 epochs, which is the number used in the paper.
@@
We show the effect of using different metrics for computing pair-wise patch similarity in 16(c). As the table shows, Cosine similarity is consistently better than Euclidean distance. We have used cosine similarity in all our experiments.
@@
ROI-Align component can be removed if the multi-crop augmentation is omitted or by using the same Global Crop across the teacher and student branches. We ablate the effect of removing ROI align in 16(d). As shown by the results, this removal negatively affects performance due to using weaker self-supervised supervision through the augmentations.
Table 16: Ablations of the key parameters of our method. We evaluate the models by training a linear layer on top of the frozen representations (Lin.) or using the in-context (IC) evaluation of Balazevic et al. (2023) using the validation images for PascalVOC12 and ADE20k. Note, that the In-Context Learning (IC) is done on the $1/128$ fraction of each dataset used.